# Deformation quantization and perverse sheaves

### v1

@@ Abstract.
Kashiwara, Polesello, Schapira and D’Agnolo defined canonical deformation quantizations of a holomorphic symplectic manifold and a holomorphic Lagrangian submanifold equipped with an orientation data. The goal of this paper is to describe the RHom complex between two such deformation quantization modules associated to a pair of Lagrangian submanifolds in terms of the derived geometry of the Lagrangian intersection. Namely, we identify the RHom complex with the DT sheaf associated to the d-critical structure on the Lagrangian intersection.
@@ Introduction
Deformation quantization modules
Let $Y$ be a holomorphic contact manifold. Kashiwara [Kas96] has defined a filtered $\mathbf{C}$-algebroid $\mathsf{E}_{Y}$ on $Y$. It provides a deformation quantization of $Y$ in the sense that $\operatorname{gr}\mathsf{E}_{Y}$ is isomorphic to $\oplus_{m\in\mathbf{Z}}\mathcal{O}_{Y}(m)$ as a sheaf of Poisson algebras. Locally $Y$ is isomorphic to an open subset of the projectivized cotangent bundle $\mathrm{P}^{*}X$ of a complex manifold $X$ in which case $\mathsf{E}_{Y}$ reduces to the algebra of microdifferential operators introduced by Sato, Kawai and Kashiwara [SKK73]. Given a holomorphic Legendrian submanifold $\Lambda\subset Y$ there is, moreover, a classification of regular holonomic $\mathsf{E}_{Y}$-modules supported on $\Lambda$ (see [Kas96, Proposition 4] and its corrected version [DS07, Corollary 6.4]) in terms of twisted local systems on the $\mathbf{C}^{\times}$-torsor $\tilde{\Lambda}\rightarrow\Lambda$ constructed from the symplectization $\tilde{Y}\rightarrow Y$.
These results were generalized from the contact to the symplectic setting in [PS04, DS07]. Namely, given a holomorphic symplectic manifold $S$ Polesello and Schapira [PS04] have defined a $\mathbf{C}(\!(\hbar)\!)$-algebroid $\mathsf{W}_{S}$ together with a $\mathbf{C}[\![\hbar]\!]$-lattice $\mathsf{W}_{S}(0)\subset\mathsf{W}_{S}$. It again provides a deformation quantization of $S$ in the sense that $\mathsf{W}_{S}(0)/\hbar$ is isomorphic to $\mathcal{O}_{S}$ as a sheaf of Poisson algebras. The construction of $\mathsf{W}_{S}$ is based on the following property: given a contactification $\rho\colon Y\rightarrow S$ of the symplectic manifold $S$, there is an embedding $\mathsf{W}_{S}\subset\rho_{*}\mathsf{E}_{Y}$ which identifies $\mathsf{W}_{S}$ with microdifferential operators $\mathsf{E}_{Y}$ on $Y$ commuting with $\partial_{t}$ (the translation vector field along the fibers of $\rho$). Under this correspondence $\hbar=\partial_{t}^{-1}$. Given a holomorphic Lagrangian submanifold $L\subset S$ equipped with a twisted local system and a number $\lambda\in\mathbf{C}/\mathbf{Z}$ there is a regular holonomic $\mathsf{W}_{S}$-module supported on $L$. In a contactification $\rho\colon Y\rightarrow S$ with a lift of $L\subset S$ to a holomorphic Legendrian $\Lambda\subset Y$ it corresponds to a regular holonomic $\mathsf{E}_{Y}$-module supported on $\Lambda$ associated to a twisted local system on $\tilde{\Lambda}\cong L\times\mathbf{C}^{\times}$ with the monodromy $\exp(2\pi i\lambda)$ along the $\mathbf{C}^{\times}$ direction. In particular, if $L$ is equipped with a square root $K_{L}^{1/2}$ of the canonical bundle, there is a simple $\mathsf{W}_{S}$-module $\mathsf{M}^{\lambda}_{L}$ supported on $L$.
Kashiwara and Schapira [KS12] have developed a general theory of deformation quantization modules. In particular, they have shown (see [KS12, Theorem 7.2.3]) that for any regular holonomic $\mathsf{W}_{S}$-modules $\mathcal{M}_{1},\mathcal{M}_{2}$ the complexes of sheaves $\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathcal{M}_{1},\mathcal{M}_{2})[\dim S/2]$ and $\mathcal{M}_{1}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}_{2}$ are perverse sheaves of $\mathbf{C}(\!(\hbar)\!)$-vector spaces. The main goal of this paper is to describe these perverse sheaves in terms of the derived geometry of the Lagrangian intersection $L\cap M$ of the supports of $\mathcal{M}_{1}$ and $\mathcal{M}_{2}$ when both $L$ and $M$ are smooth holomorphic Lagrangian submanifolds of $S$.
Microlocal sheaves and Fukaya categories
In this section we outline connections between the theory of deformation quantization modules and Fukaya categories, on the one hand, and microlocal sheaves, on the other hand, to place our results in a broader context.
As the first goal let us describe the relevant category of deformation quantization modules associated to a holomorphic symplectic manifold $S$ (we review this construction in Section 4). Fix a (possibly singular) complex Lagrangian subvariety $L\subset S$. There is a canonical contactification $\rho\colon Y\rightarrow S$ of a neighborhood of $L$ with a lift of $L\subset S$ to a Legendrian subvariety $\Lambda\subset Y$ (see 1.12). As in [DK11] we define the abelian stack $\mathrm{RH}_{S,L}$ (i.e. a sheaf of $\mathbf{C}$-linear abelian categories) of regular holonomic DQ modules on $S$ supported on $L$ to be the stack of regular holonomic $\mathsf{E}_{Y}$-modules supported on $\Lambda$, i.e.
$\mathrm{RH}_{S,L}=\rho_{*}\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y}).$
Lagrangian subvarieties of $S$ form a filtered system and we define $\mathrm{RH}_{S}$ as the filtered colimit
$\mathrm{RH}_{S}=\operatorname*{colim}_{L}\mathrm{RH}_{S,L}.$
In particular, if $L\subset S$ is a smooth Lagrangian submanifold equipped with a square root line bundle $K_{L}^{1/2}$ (orientation data) and $\lambda\in\mathbf{C}/\mathbf{Z}$ we have the objects $\mathsf{M}^{\lambda}_{L}\in\mathrm{RH}_{S,L}$ mentioned before.
Let us now describe a topological avatar of the stack $\mathrm{RH}_{S}$. Let $W$ be an exact holomorphic symplectic manifold. The authors of [NS20, Côt+22] have introduced a sheaf of stable $\infty$-categories $\mu\mathrm{sh}_{W,\mathbf{C}\text{-}c}$ on $W$ by globalizing the microlocal theory of sheaves of Kashiwara and Schapira [KS94]. For a holomorphic contact manifold $Y$ one then defines the stack $\mathbb{P}\mu\mathrm{sh}_{Y,\mathbf{C}\text{-}c}=\gamma_{*}\mu\mathrm{sh}_{\tilde{Y},\mathbf{C}\text{-}c}$, where $\gamma\colon\tilde{Y}\rightarrow Y$ is the symplectization. One has
$\mathbb{P}\mu\mathrm{sh}_{Y,\mathbf{C}\text{-}c}=\operatorname*{colim}_{\Lambda}\mathbb{P}\mu\mathrm{sh}_{\Lambda},$
where $\mathbb{P}\mu\mathrm{sh}_{\Lambda}\subset\mathbb{P}\mu\mathrm{sh}_{Y,\mathbf{C}\text{-}c}$ is the subcategory of objects supported on a Legendrian subvariety $\Lambda\subset Y$. Moreover, the authors of [Côt+22] construct a $t$-structure on $\mathbb{P}\mu\mathrm{sh}_{Y,\mathbf{C}\text{-}c}$ (respectively, $\mathbb{P}\mu\mathrm{sh}_{\Lambda}$) whose heart is the stack $\mu\mathrm{Perv}_{Y}$ (respectively, $\mu\mathrm{Perv}_{Y,\Lambda}$) of microlocal perverse sheaves as in [And96, Was04].
For $Y=\mathrm{P}^{*}X$ a projectivized cotangent bundle and $\Lambda\subset Y$ a fixed Legendrian subvariety Waschkies [Was05] has defined a microlocal Riemann–Hilbert correspondence which identifies
$\mu\mathrm{Perv}_{Y}\cong\mathrm{Mod}_{\mathrm{rh}}(\mathsf{E}_{Y}),\qquad\mu\mathrm{Perv}_{Y,\Lambda}\cong\mathrm{Mod}_{\Lambda,\mathrm{rh}}(\mathsf{E}_{Y}),$
which explains a connection between the stack $\mathrm{RH}_{S}$ and the stacks of microlocal sheaves.
If the exact holomorphic symplectic manifold $\tilde{Y}$ is Weinstein, then Ganatra, Pardon and Shende [GPS18] identify $\mu\mathrm{sh}_{\tilde{Y}}$ with the ind-completion of the wrapped Fukaya category of $\tilde{Y}$. Thus, we have three independent ways to approach the stack $\mathrm{RH}_{S}$: in terms of deformation quantization modules, in terms of microlocal sheaf theory and in terms of Fukaya categories.
D-critical loci and DQ modules
The relevant derived structure on a Lagrangian intersection that we will need is the structure of a d-critical locus introduced in [Joy15]. We recall the definition in Section 6.2, but let us simply mention that a d-critical locus is locally isomorphic to the critical locus $\mathrm{Crit}(f)$ of a holomorphic function $f\colon U\rightarrow\mathbf{C}$ on a complex manifold $U$ as a complex-analytic space. Any d-critical locus has an orientation $\mathbf{Z}/2$-gerbe whose trivialization is the choice of an orientation data. The key result of [Bra+15] is a construction of a canonical perverse sheaf $\phi_{X}$ (of $\mathbf{C}$-vector spaces), which we call the DT sheaf, on an oriented d-critical locus globalizing the sheaf of vanishing cycles $\phi_{f}$. (More canonically, we can think of $\phi_{X}$ as a sheaf twisted by the orientation gerbe similar to the fundamental class of a closed manifold being a homology class twisted by the orientation local system.)
As explained in [Bus14], if $S$ is a holomorphic symplectic manifold and $L,M\subset S$ are holomorphic Lagrangian submanifolds, the Lagrangian intersection $L\cap M$ has a canonical structure of a d-critical locus. Moreover, if $L$ and $M$ are equipped with square roots $K_{L}^{1/2},K_{M}^{1/2}$ of the canonical bundles, then the d-critical locus $L\cap M$ is oriented; in particular, there is the DT sheaf $\phi_{L\cap M}$ on the Lagrangian intersection. We may then state the main result (7.1) as follows.
Theorem A.
@@
For any $\lambda,\mu\in\mathbf{C}$ there is an isomorphism of $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaves
@@
$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathsf{M}^{\mu}_{M}\cong\phi_{L\cap M}\otimes_{\mathbf{C}}\mathbf{C}(\!(\hbar)\!).$
When the Lagrangians $L,M$ intersect cleanly, the right-hand side reduces to an explicit $\mathbf{Z}/2$-torsor described in Section 7.4. This fact should be compared with a similar computation of Lagrangian Floer homology of clean intersections [Poź99, Sch16].
In fact, we prove a strengthening of the above statement which explains the origin of the $\hbar$ parameter. For this, recall that the DT sheaf $\phi_{L\cap M}$ carries a canonical monodromy automorphism. The Riemann–Hilbert correspondence identifies perverse sheaves $P$ of $\mathbf{C}$-vector spaces with an automorphism and perverse sheaves $\mathrm{RH}^{-1}(P)$ of $\mathbf{C}(\!(\hbar)\!)$-vector spaces equipped with a connection in the $\hbar$-direction with regular singularities (see 2.11). We call perverse sheaves of $\mathbf{C}(\!(\hbar)\!)$-vector spaces equipped with connections in the $\hbar$-direction differential perverse sheaves. We also equip the derived tensor product $\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathsf{M}^{\mu}_{M}$ with a natural connection in the $\hbar$-direction coming from the fact that the $\mathsf{W}_{S}$-modules $\mathsf{M}^{\lambda}_{L},\mathsf{M}^{\mu}_{M}$ locally come from $\mathsf{E}_{Y}$-modules on a contactification. The stronger statement of the main result (7.1) is then as follows.
Theorem B.
$\mathsf{M}^{\lambda}_{L}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathsf{M}^{\mu}_{M}\cong\mathrm{RH}^{-1}(\phi_{L\cap M}\otimes\mathbf{C}_{\mu+\lambda}),$
@@
One motivation for this statement is a proposal [Bra+15, Remark 6.15] for a Fukaya-like category associated to a holomorphic symplectic manifold $S$:
@@
  • Composition of Homs is given by a conjectural functoriality of the DT sheaf, see [JS19, Conjecture 1.1].
We refer to [SV19] and [DR22, Section B.2] for a connection between this category and Lagrangian Floer homology. 7.1 shows that the stack $\mathrm{RH}_{S}$ provides such a category. Indeed, there is an isomorphism $\mathcal{M}_{1}\otimes^{\mathbb{L}}_{\mathsf{W}_{S}}\mathcal{M}_{2}\cong\mathbb{R}\mathcal{H}om_{\mathsf{W}_{S}}(\mathbb{D}(\mathcal{M}_{1}),\mathcal{M}_{2})[\dim S/2]$ valid for any holonomic $\mathsf{W}_{S}$-modules $\mathcal{M}_{1},\mathcal{M}_{2}$, where the duality functor is $\mathbb{D}(\mathcal{M}_{1})=\mathbb{R}\mathcal{H}om(\mathcal{M}_{1},\mathsf{W}_{S})[\dim S/2]$. Moreover, the modules $\mathsf{M}^{\lambda}_{L}$ are self-dual.
The format of the statement of 7.1 is inspired by a related result [SS14] of Sabbah and Saito (see also [Sab10]) relating the twisted de Rham complex
and the sheaf of vanishing cycles $\phi_{f}$. (Note that we restrict to the zero set of $f$ to simplify the presentation.) Namely, there is a natural connection on $\mathrm{DR}_{f}$ in the $\hbar$ direction given by $\partial_{\hbar}-\hbar^{-2}f$, which equips $\mathrm{DR}_{f}$ with the structure of a differential perverse sheaf. Sabbah and Saito [SS14] identify the cohomology sheaves $\mathcal{H}^{k}\mathrm{DR}_{f}$ and $\mathcal{H}^{k}\mathrm{RH}^{-1}(\phi_{f})$. We show that their isomorphism lifts to an isomorphism of differential perverse sheaves (see 5.11).
Theorem C.
A closely related result has recently been shown by Schefers in [Sch23, Proposition 9.20] using different methods.
@@
Another motivation for this work is a conjectural relationship between skein modules and the DT sheaf formulated in [GJS23]. Let us recall the statement. Let $(M,x)$ be a connected closed oriented 3-manifold and $G$ a reductive algebraic group equipped with some extra data described in [GJS23, Section 1.3]. Then one can define the skein module $\mathrm{Sk}^{\mathrm{gen}}_{G}(M)$ of $M$ for generic quantum parameters, which is a $\mathbf{C}(q^{1/d})$-vector space for some integer $d$. For instance, for $G=\mathrm{SL}_{2}$ we obtain the Kauffman bracket skein module, which is a $\mathbf{C}(A)$-vector space for $A=q^{1/2}$. The representation variety
@@
carries a natural structure of a d-critical locus. In the case $G=\mathrm{SL}_{2}$ Abouzaid and Manolescu [AM20] denote the hypercohomology of the DT sheaf $\phi_{R_{G}(M)}$ on $R_{G}(M)$ by $\mathrm{HP}^{\bullet}_{\sharp}(M)$, the framed complexified instanton Floer homology of $M$. The conjectural relationship between the two objects is as follows.
Conjecture D.
@@
We refer to [DKS23] for some checks of this conjecture for the Kauffman bracket skein module.
The above conjecture is related to A as follows. Choose a Heegaard splitting $M=N_{1}\cup_{\Sigma}N_{2}$ of $M$ as a union of two handlebodies $N_{1},N_{2}$ along the common boundary $\Sigma$. In this case we have the following:
  • The representation variety $R_{G}(M)$ becomes an intersection $L_{1}\cap L_{2}$ of two complex algebraic Lagrangian submanifolds $L_{1},L_{2}$ (representation varieties of the handlebodies $N_{1},N_{2}$) in a complex symplectic variety $S$ (an open subset of the representation variety of the open surface $\Sigma^{*}=\Sigma\setminus D$). We refer to [GJS23, Section 5.3] for details.
@@
Thus, in order to prove D using A, one must in particular relate the tensor product of the algebraic deformation quantization provided by skein theory to the corresponding tensor product of the canonical analytic deformation quantization modules. We plan to address this problem in future work.
@@
  • For a complex manifold $X$ and a ring $R$ we denote by $\mathbf{D}(X;R)$ the derived category of sheaves of complexes of $R$-modules.
  • $\mathbf{D}_{c}(X;R)\subset\mathbf{D}(X;R)$ denotes the derived category of constructible sheaves of $R$-modules on $X$.

### v2

@@ Abstract.
Kashiwara, Polesello, Schapira and D’Agnolo defined canonical deformation quantizations of a holomorphic symplectic manifold and a holomorphic Lagrangian submanifold equipped with an orientation data. The goal of this paper is to use deformation quantization modules to construct a Fukaya-like category of holomorphic Lagrangians, resolving a conjecture of Joyce. Our main result describes the RHom complex between two such deformation quantization modules associated to a pair of Lagrangian submanifolds in terms of the derived geometry of the Lagrangian intersection. Namely, we identify the RHom complex with the DT sheaf associated to the d-critical structure on the Lagrangian intersection. Via a Riemann–Hilbert correspondence this describes the RHom complex in the category of microsheaves of sheaf quantizations of conic holomorphic Lagrangians.
@@ Introduction
Suppose we are given a holomorphic symplectic manifold $S$ and a pair of Lagrangian submanifolds $L,M$ each equipped with a choice of square root of the canonical bundle. Associated to this data we have the following two perverse sheaves on the intersection $L\cap M$:
  • By [Bussi, BBDJS], the intersection $L\cap M$ admits the structure of a $d$-critical locus, and thus carries a canonical $\mathbf{C}$-linear perverse sheaf $\phi_{L\cap M}$ (called the DT sheaf), equipped with a canonical automorphism $T_{L\cap M}$ (called the monodromy), obtained by gluing together vanishing cycle sheaves on local charts.
  • By [PoleselloSchapira, DAgnoloSchapira, KashiwaraSchapira], there is a canonical $\mathbf{C}(\!(\hbar)\!)$-algebroid $\widehat{\mathsf{W}}_{S}$ together with modules $\mathsf{M}_{L}$, $\mathsf{M}_{M}$, and
    $\mathsf{M}_{L}\otimes^{L}_{\widehat{\mathsf{W}}_{S}}\mathsf{M}_{M}\cong\mathbb{R}\mathcal{H}om_{\widehat{\mathsf{W}}_{S}}(\mathsf{M}_{L},\mathsf{M}_{M})[\dim S]$
    is a $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaf supported on $L\cap M$.
Our primary goal of this paper is to compare these two perverse sheaves. To put the two sides on the same footing, we use a version of the Riemann-Hilbert correspondence to reinterpret the pair $(\phi_{L\cap M},T_{L\cap M})$ as a differential perverse sheaf $\mathrm{RH}^{-1}(\phi_{L\cap M},T_{L\cap M})$ (informally, a $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaf equipped with a regular $\hbar$-connection).
Theorem A (7.1).
There is a natural differential structure on the perverse sheaf $\mathsf{M}_{L}\otimes^{L}_{\widehat{\mathsf{W}}_{S}}\mathsf{M}_{M}$, and an isomorphism of differential perverse sheaves
$\mathsf{M}_{L}\otimes^{\mathbb{L}}_{\widehat{\mathsf{W}}_{S}}\mathsf{M}_{M}\cong\mathrm{RH}^{-1}(\phi_{L\cap M},T_{L\cap M}).$
@@
It follows from A that there is an isomorphism of $\mathbf{C}(\!(\hbar)\!)$-linear perverse sheaves
@@
$\mathbb{R}\mathcal{H}om_{\widehat{\mathsf{W}}_{S}}(\mathsf{M}_{L},\mathsf{M}_{M})[\dim S]\cong\phi_{L\cap M}\otimes_{\mathbf{C}}\mathbf{C}(\!(\hbar)\!)$
While such a comparison has been predicted for some time in the community (see [Bussi, Section 4]), this is the first proof of this result known to the authors and also the first statement that properly accounts for the monodromy and discrepancy of coefficient fields. As we explain below, the extra structure corresponding to the monodromy appears naturally in our approach and is a necessary ingredient in our proof of any such comparison.
A has a number of consequences and potential applications, which we explore in the rest of this introduction: to skein modules, microlocal sheaves and Fukaya categories.
A holomorphic Fukaya category
@@
One source of motivation for our work is a proposal [BBDJS, Remark 6.15] for a Fukaya-like category associated to a holomorphic symplectic manifold $S$:
@@
  • Composition of Homs is given by a conjectural functoriality of the DT sheaf, see [JoyceSafronov, Conjecture 1.1].
Our results provide the first construction of such a category (at least over $\mathbf{C}$). Let $\mathrm{Rep}(\mathbf{Z})$ be the $\infty$-category of complexes of $\mathbf{C}$-vector spaces equipped with an automorphism.
Theorem B (8.5).
Let $S$ be a holomorphic symplectic manifold. There is a sheaf $\mathbf{D}\mathsf{RH}_{S}$ of $\mathrm{Rep}(\mathbf{Z})$-linear $\infty$-categories on $S$, equipped with a $t$-structure, which satisfies the following properties:
  • For each Lagrangian submanifold $L$ of $S$ equipped with orientation data, there is a canonical object $\mathsf{M}_{L}$.
  • Given a pair of oriented Lagrangian submanifolds $L,M$, there is an isomorphism
    $\underline{\mathcal{H}om}_{\mathbf{D}\mathsf{RH}_{S}}(\mathsf{M}_{M},\mathsf{M}_{L})[\dim S/2]\cong(\phi_{L\cap M},T_{L\cap M})$
    of perverse sheaves equipped with an automorphism.
@@
Another motivation for this work is a conjectural relationship between skein modules and the DT sheaf formulated in [GJS]. Let us recall the statement. Let $M$ be a connected closed oriented 3-manifold with a basepoint $x\in M$ and $G$ a reductive algebraic group equipped with some extra data described in [GJS, Section 1.3]. Then one can define the skein module $\mathrm{Sk}^{\mathrm{gen}}_{G}(M)$ of $M$ for generic quantum parameters, which is a $\mathbf{C}(q^{1/d})$-vector space for some integer $d$. For instance, for $G=\mathrm{SL}_{2}$ we obtain the Kauffman bracket skein module, which is a $\mathbf{C}(A)$-vector space for $A=q^{1/2}$. The representation variety
@@
carries a natural structure of a d-critical locus. In the case $G=\mathrm{SL}_{2}$ Abouzaid and Manolescu [AbouzaidManolescu] denote the hypercohomology of the DT sheaf $\phi_{R_{G}(M)}$ on $R_{G}(M)$ by $\mathrm{HP}^{\bullet}_{\sharp}(M)$, the framed complexified instanton Floer homology of $M$. The conjectural relationship between the two objects is as follows.
Conjecture C.
@@
We refer to [DKS] for some checks of this conjecture for the Kauffman bracket skein module. The above conjecture is related to A as follows. Choose a Heegaard splitting $M=N_{1}\cup_{\Sigma}N_{2}$ of $M$ as a union of two handlebodies $N_{1},N_{2}$ along the common boundary $\Sigma$. In this case we have the following:
  • The representation variety $R_{G}(M)$ becomes an intersection $L_{1}\cap L_{2}$ of two complex algebraic Lagrangian submanifolds $L_{1},L_{2}$ (representation varieties of the handlebodies $N_{1},N_{2}$) in a complex symplectic variety $S$ (an open subset of the representation variety of the open surface $\Sigma^{*}=\Sigma\setminus D$). We refer to [GJS, Section 5.3] for details.
@@
Thus, in order to prove C using A, one must in particular relate the tensor product of the algebraic deformation quantization modules provided by skein theory to the corresponding tensor product of the canonical analytic deformation quantization modules. We plan to address this problem in future work.
Vanishing cycles and the twisted de Rham complex
Consider the case when $S=\mathrm{T}^{\ast}X$ is the cotangent bundle of a complex manifold $X$, $L=\Gamma_{0}$ is the zero-section and $M=\Gamma_{df}$ is the graph of $df$, where $f\colon X\to\mathbf{C}$ is a holomorphic function. This can be considered as a local model for the setup of A, in the sense that every holomorphic Lagrangian intersection is locally isomorphic to such an intersection. In this case, the DT sheaf $\phi_{\Gamma_{0}\cap\Gamma_{df}}$ is identified with the sheaf of vanishing cycles $\phi_{f}$.
The format of the statement of A is inspired by a related result [SabbahSaito] of Sabbah and Saito (see also [Sabbah]) relating the twisted de Rham complex
and the sheaf of vanishing cycles $\phi_{f}$. (Note that we restrict to the zero set of $f$ to simplify the presentation.) Namely, there is a natural connection on $\mathrm{DR}_{f}$ in the $\hbar$ direction given by $\partial_{\hbar}-\hbar^{-2}f$, which equips $\mathrm{DR}_{f}$ with the structure of a differential perverse sheaf. Sabbah and Saito [SabbahSaito] identify the cohomology sheaves $\mathcal{H}^{k}\mathrm{DR}_{f}$ and $\mathcal{H}^{k}\mathrm{RH}^{-1}(\phi_{f})$. We show that their isomorphism lifts to an isomorphism of differential perverse sheaves.
Theorem D (5.11).
A closely related result has recently been shown by Schefers in [Schefers, Proposition 9.20] using different methods.
Contactification and symplectization
The results and techniques in this paper are rooted in the tight interplay between symplectic and contact geometry in the holomorphic setting. Understanding this interplay also leads to a conceptual explanation for the structure appearing in the holomorphic Fukaya category of B.
Given a pair $(S,L)$ of a symplectic manifold together with a Lagrangian subvariety (for example, a finite union of Lagrangian submanifolds), a contactification of $(S,L)$ is a pair $(Y,\rho,\Lambda)$ where $\rho:Y\to S$ is a principal $\mathbf{C}$-bundle with connection whose curvature is the symplectic form, $\Lambda$ is a Legendrian subvariety of $Y$ (with respect to the canonical contact structure), and $\rho|_{\Lambda}$ is a homeomorphism onto $L$, analytic on the smooth part of $\Lambda$.
Given an arbitrary contact manifold $Y$, there is a canonical $\mathbf{C}^{\times}$-principal bundle $\gamma:\widetilde{Y}\to Y$ called the symplectization. In the case when $\rho:Y\to S$ is the contactification of a symplectic manifold $S$, the symplectization is canonically trivial: $\widetilde{Y}=Y\times\mathbf{C}^{\times}$. Moreover, if $L$ is a Lagrangian subvariety of $S$, and $(Y,\Lambda)$ a contactification of $(S,L)$, then $\widetilde{\Lambda}=\gamma^{-1}(\Lambda)\cong\Lambda\times\mathbf{C}^{\times}$ is an exact Lagrangian subvariety in $\widetilde{Y}$.
Note that in our main results we do not assume that our symplectic manifold $S$ is exact. In particular, there may not exist a global contactification $\rho:Y\to S$. Nevertheless, we have the following fundamental result (see [DAgnoloKashiwara, Proposition 2.5.3]).
Proposition (Unique Contactification Lemma, 1.12).
Any pair $(S,L)$ of a symplectic manifold and Lagrangian subvariety admits a contactification $(Y,\rho,\Lambda)$ over a neighborhood of $L$ in $S$. Moreover, this data is unique up to a suitable notion of isomorphism.
An important consequence of this result is that, given an arbitrary Lagrangian intersection $L\cap M$ in a symplectic manifold $S$, there is a canonical Legendrian intersection $\Lambda_{L}\cap\Lambda_{M}$ in a contactification $Y$ of $S$ (indeed, consider the unique contactification of $(S,L\cup M)$). In particular, $(L\cap M)\times\mathbf{C}^{\times}$ is homeomorphic to the intersection $\widetilde{\Lambda}_{L}\cap\widetilde{\Lambda}_{M}$ of $\mathbf{C}^{\times}$-invariant (exact) Lagrangians in the symplectization $\widetilde{Y}$. As explained in Section 8, the various perverse sheaves associated to the Lagrangian intersection $L\cap M$ in $S$ can be realized naturally as perverse sheaves on $\widetilde{\Lambda}_{L}\cap\widetilde{\Lambda}_{M}\cong(L\cap M)\times\mathbf{C}^{\times}$ which are lisse along the $\mathbf{C}^{\times}$-direction.
We are now in a position to recall the construction, due to D’Agnolo–Kashiwara [DAgnoloKashiwara], of the category $\mathsf{RH}_{S}$ on an arbitrary symplectic manifold $S$. Suppose $L\subseteq S$ is a Lagrangian subvariety and let $(Y,\Lambda)$ be a contactification in a neighborhood of $L$ (which exists by the unique contactification lemma 1.12). Kashiwara [KashiwaraContact] has defined a certain filtered $\mathbf{C}$-algebroid $\mathsf{E}_{Y}$ on a general contact manifold $Y$, locally modelled on the algebra of microdifferential operators $\mathcal{E}_{X}$ on the projectivized cotangent bundle $\mathrm{P}^{\ast}X$ of a complex manifold $X$. We denote by $\mathsf{RH}_{S,L}$ the category of regular holonomic $\mathsf{E}_{Y}$-modules along $\Lambda$ (understood as a sheaf of categories on $L$ by pushing forward along the homeomorphism $\rho|_{\Lambda}$). We then define $\mathsf{RH}_{S}$ by taking the filtered colimit of $\mathsf{RH}_{S,L}$ along the diagram of all Lagrangian subvarieties $L$ in $S$. The uniqueness part of the unique contactification lemma tells us that this construction is well-defined.
Microlocal sheaves
For an exact holomorphic symplectic manifold $W$ the authors of [NadlerShende, CKNS1] have introduced a sheaf of stable $\infty$-categories $\mu\mathrm{sh}_{W}$ on $W$ of microsheaves by globalizing the microlocal theory of sheaves of Kashiwara and Schapira [KashiwaraSchapiraSheaves]. Moreover, for a holomorphic contact manifold $Y$ the authors of [CKNS1] have defined a sheaf of abelian categories $\mathbb{P}\mathrm{erv}_{Y}\subset\gamma_{*}\mu\mathrm{sh}_{\tilde{Y}}$ of perverse microsheaves following [Andronikof2, Waschkies1]. Perverse microsheaves relate to deformation quantization modules via a microlocal Riemann–Hilbert correspondence [Andronikof3, Waschkies2, CKNS2]: namely, there is an equivalence of sheaves of abelian categories
$\mu\mathrm{RH}\colon\mathbb{P}\mathrm{erv}_{Y}\xrightarrow{\sim}\mathrm{Mod}_{\mathrm{rh}}(\mathsf{E}_{Y}).$
Given a holomorphic Legendrian submanifold $\Lambda\subset Y$ equipped with a square root line bundle $K_{\Lambda}^{1/2}$ there is a perverse microsheaf $\mathcal{M}_{\Lambda}\in\mathbb{P}\mathrm{erv}_{Y}$ supported on $\Lambda$ which provides a sheaf quantization of $\Lambda$ in the sense that its microlocalization to $\Lambda$ is a trivial (twisted) local system of rank $1$. Equivalently, $\mu\mathrm{RH}(\mathcal{M}_{\Lambda})=\mathsf{M}_{\Lambda}$ is a simple $\mathsf{E}_{Y}$-module along $\Lambda$. B allows us to compute Hom complexes between two such sheaf quantizations.
Theorem E (8.8).
Let $Y$ be a holomorphic contact manifold and $\Lambda_{1},\Lambda_{2}\subset Y$ two holomorphic Legendrian submanifolds equipped with square root line bundles $K_{\Lambda_{1}}^{1/2},K_{\Lambda_{2}}^{1/2}$. Let $\tilde{\Lambda}_{1},\tilde{\Lambda}_{2}$ be the corresponding homogeneous Lagrangian submanifolds of the symplectization $\tilde{Y}\rightarrow Y$. There is an isomorphism of complexes of sheaves
$\mathcal{H}om_{\mu\mathrm{sh}_{\tilde{Y}}}(\mathcal{M}_{\Lambda_{1}},\mathcal{M}_{\Lambda_{2}})[\dim\tilde{Y}/2]\cong\phi_{\tilde{\Lambda}_{1}\cap\tilde{\Lambda}_{2}}$
on $\tilde{Y}$.
Fukaya categories
Consider, as before, a holomorphic contact manifold $Y$ with a symplectization $\tilde{Y}\rightarrow Y$. In particular, $\tilde{Y}$ is a real exact symplectic manifold. If we assume that it is Weinstein, then Ganatra, Pardon and Shende [GPS] identify $\mu\mathrm{sh}_{\tilde{Y}}$ with the ind-completion of the wrapped Fukaya category of $\tilde{Y}$. In particular, in this case the left-hand side of E may be rewritten in terms of FLoer homology. Via the microlocal Riemann–Hilbert correspondence, this relates the sheaf of $\infty$-categories $\mathbf{D}\mathsf{RH}_{S}$ on a (non-exact) holomorphic symplectic manifold $S$ constructed in B to Floer theory in the symplectization $\tilde{Y}$ of a local contactification $Y\rightarrow S$.
The relationship between the sheaf of $\infty$-categories $\mathbf{D}\mathsf{RH}_{S}$ and Lagrangian Floer theory in $S$ is more subtle, see the discussion in [DoanRezchikov, Section B.2] and [KontsevichSoibelman, Section 6.3]. In this direction, let us mention the locality results for holomorphic Floer theory established by Solomon and Verbitsky in [SolomonVerbitsky] as well as a relationship between Lagrangian Floer theory in $S$ and that on the symplectization of a real contactification established by Kuwagaki, Petr and Shende in [KuwagakiPetrShende].
@@
  • For a complex manifold $X$ and a ring $R$ we denote by $\mathbf{D}(X;R)$ the derived $\infty$-category of sheaves of complexes of $R$-modules.
  • $\mathbf{D}_{c}(X;R)\subset\mathbf{D}(X;R)$ denotes the subcategory of constructible complexes of $R$-modules on $X$.
@@
  • A stack of categories is the same as a sheaf of categories. When there is no confusion, we will denote its global sections by the same letter.